A pipe ca fill a pool in 9 hours. The drain can empty the pool in 12 hours. I f the drain pipe is left open, how long will it take to fill the pool?
Answer by dolly(163)

Time taken to fill the pool = 9 hrs
Volume filled in 1 hr = 1/9 part of the pool
Time taken to drain = 12 hrs.
Volume drained in 1 hr = 1/12 part
When drain is left open, volume filled in 1 hr = 1/9 - 1/12
= (4-3)/36
= 1/36
Thus time taken to fill the pool = 36 hrs when the drain is open